A lacrosse ball is hurled straight up, leaving the head of the stick at a height of 2.00 m above the ground level. The ball passes by a 1.25-m-high window in a time of 0.350 s as it heads upward. 1) Calculate the initial speed of the ball if it is 13.0 m from the ground to the bottom of the window. Ignore the effects of air resistance. (Express your answer to three significant figures.) m Submit
Answer #1

Solution)

We know, s=ut + 0.5 at^2

1.25 m=v*0.350 + 0.5 *9.8*0.350^2

So, v= 1.856 m/s

Now, we know

V^2 - u^2=2as

U= sqrt( v^2 + 2gy)

U= sqrt( 1.856^2 + 2*9.8*11) =14.80 m/s (Ans)

======